i need to know how to find the equation of the line containing the given pair of points. express your in answer in the form of a x=a,y=mx+b. the problum is (-1,-1).and (8,8) and the other part is. what is and an equation of the line y =
Answer by rfer(16322)

slope m= (8--1)/(8--1)
m=9/9
m=1
---------------
y=mx+b
8=8+b
b=0
----------------
y=x